OpenStudy (anonymous):

Three cards labeled A, B, and C are placed face down on a table. A number cube has faces labeled with the numbers 1 through 6. What is the probability of turning over a card labeled B and rolling a number greater than 2 on the number cube?

OpenStudy (anonymous):

A.2/9 B.5/18 C.1/3 D. 1/2

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

i think a

OpenStudy (anonymous):

you have a 1/3 chance to pick b and 2/3 to roll a number higher then 2

OpenStudy (welshfella):

first probability = 1/3 and second = 2/3 multiply them

OpenStudy (anonymous):

2/9 @welshfella

OpenStudy (anonymous):

it wassss corrrrrrreeeeeeecccctttttt